OpenStudy (anonymous):

Differentiate y= (1-x^2) (1+4x)^3 with respect to x.

OpenStudy (sriram):

using product rule, (1-x^2)*3*(1+4x)^2*4 +(1+4x)^3*(-2x)

OpenStudy (anonymous):

I learn the differentiate first term keep seceond term then keep first term differentiate second term method

OpenStudy (anonymous):

\(y'=(1-x^2)(3(4)(1+4x)^2)+(1+4x)^3(-2x)\) \(=12(1-x^2)(1+4x)^2-2x(1+4x)^3=(1+4x)^2(12-12x^2-2x-8x^2)\) \(=(1+4x)^2(-20x^2-2x+12)\).

OpenStudy (anonymous):

Anwara what is the final answer?

OpenStudy (anonymous):

Well, you can take this as a final answer, I mean the last line. You can also factorize the second parenthesis by taking \(2\) as a common factor.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

anwar you are awesomE =d

OpenStudy (anonymous):

I know, but thanks :D